OpenStudy (anonymous):

Determine the points of discontinuity. State the type of discontinuity (removable, jump, infinite, or none of these) and whether the function is left- or right-continuous. 1. f(x) = lxl 2. g(t) = 1/(t^2-1) 3. f(x) = { cos (1/x) x does not = 0 { 1 x = 0 4. f(x) = ln l x-4 l

OpenStudy (anonymous):

absolute value is continuous

OpenStudy (anonymous):

Thank you that is what I thought for the that one! I also believe the answer to the second one is that there is an infinite discontinuity at t =1 and t=-1. But I am not positive. Is the answer to the fourth that there is an infinite discontinuity when x<4 and it is right continuous? The third one is giving me the most confusion.

OpenStudy (anonymous):

yes for number 2

OpenStudy (anonymous):

and number 4 since the log of zero is undefined, and it goes to \(-\infty\) there

OpenStudy (anonymous):

for the third one, you have to compute the limit as \(x\to 0\) of \(\cos(\frac{1}{x})\) if that limit exists, since cosine is continuous everywhere you can remove the discontinuity, or if that limit is 1, your function is continuous everywhere

OpenStudy (anonymous):

unfortunately, as \(x\to 0^+\) we know \(\frac{1}{x}\to \infty\) and cosine will not have a limit, since it oscillated between -1 and 1 infinitely often so there is not limit, and the discontinuity is not removable

OpenStudy (anonymous):

however it not infinite limit, since it only goes between -1 and 1, not up to infinity also it is not a jump, since the function is continuous everywhere except at 0